OpenStudy (anonymous):

integrate this please. int. sqr of dv / sqr. of (2-v)

OpenStudy (anonymous):

\[\int\limits \sqrt{v} dv / \sqrt{2-v}\] this is the correct equation. sorry, i forgot sqr of v.

OpenStudy (dumbcow):

An alternative to wolframs steps: v = 2sin^2(theta) dv = 4sin(theta)cos(theta) dtheta reducing integral to \[4\int\limits_{}^{}\sin^{2}\theta d \theta \]

OpenStudy (jamesj):

Nice.

OpenStudy (anonymous):

thank you! :D